OpenStudy (angelwings996):

Algebra 2 help please... Write an equation for a horizontal translation of y = 2/x. Then write an equation for a vertical translation of y = 2/x. Identify the horizontal and vertical asymptotes of the graph of each function.

OpenStudy (mathstudent55):

A horizontal translation is usually accomplished by replacing x by x - h, where the translation is h units. For the vertical translation, you do a similar thing with y. Replace y with y - k for a transdlation of k units. Since you have y = 2/x, replacing y with y - k, you get y = 2/x + k Choose numbers of h and for k.

OpenStudy (angelwings996):

@mathstudent55 would the vertical translation be y-k=2/x ?

OpenStudy (mathstudent55):

yes. all you need to do is choose a number for k.

OpenStudy (angelwings996):

@mathstudent55 okay, so if i use 2 for k, how do I find the horizontal and vertical asymptotes?

OpenStudy (mathstudent55):

Do you know how to find the asymptotes of the original function, y = 2/x?

OpenStudy (angelwings996):

I'm not really sure how you find asymptotes period :/

OpenStudy (mathstudent55):

The asymptotes of a hyperbola of the form xy = c, where c is a non-zero number, are the x-axis and the y-axis. Since the form is xy = c, you can solve for y and get y = c/x In your case, you have y = 2/x. For the hyperbola y = 2/x, the asymptotes are the coordinate axes.

OpenStudy (angelwings996):

Oh okay, so if I was to use 2 for k then would I have to do any math to it? @mathstudent55

OpenStudy (mathstudent55):

When k = 2, the entire graph shifts 2 untis up. So the vertical asymptote remains the same, and the horizontal asymptote becomes the line x = 2, which is a horizontal line 2 units above the x-axis.